# Building imaginary-time thermal field theory with artificial neural networks

### v1

@@ III Neural Networks for Quantum Statistics
where $\Delta\tau=\beta/N_{\tau}$, the kinetic term is $K\equiv N_{\tau}\sum(\Delta x)^{3}{(\Delta\Phi)^{2}}$, and the potential term is $V\equiv N_{\tau}^{-1}\sum(\Delta x)^{3}{[(\nabla\Phi)^{2}+V(\Phi)]}$. The temperature dependence is separable with regard to the quantum fields. It is evident that once actions of a sample at two given temperatures can be estimated as
@@
$V_{k}(x)=\frac{\lambda_{k}}{4}\left(x^{2}-\frac{\mu_{k}^{2}}{2k}\right)^{2}.$ (9)
Here $\lambda_{k}$ is the coupling constant with mass dimension $[\lambda_{k}]\leavevmode\nobreak\ [m]^{5}$ and a parameter $\mu_{k}$ with mass dimension $[\mu_{k}]\leavevmode\nobreak\ [m]^{3/2}$. Suppose the particle mass parameter (in the kinetic energy term) is m, we will then work with dimensionless quantities rescaled by proper powers of m (e.g., $\mu^{2}_{k}/m^{3}\rightarrow\mu^{2}_{k}$ and $\lambda_{k}/m^{5}\rightarrow\lambda_{k}$) throughout this paper. The properties of the system at finite temperature $T$ can be described through the partition function
@@ V.1 Action Estimation
In this section we show the training and validation stages of the CANs. It will be seen that the CANs can succeed to reproduce the correct action of most configurations at different temperature without knowing the analytical form of the Lagrangian density (unsupervisely). To train the network we have prepared three ensembles at $\beta=T^{-1}=$20, 40 and 80 with the MCMC simulation. We train the CANs to estimate the probability of each sample in these three ensembles for 10,000 epochs. From Figure 2 we can be sure that network has been convergent.
Figure 2: Model loss for $beta=20,40,80$,respectively
@@
Figure 3: (a)-(c)Comparisons of action distribution by MC(analytical) and Net; (d)-(f)Comparisons of action on testing data by MC and Net; (g)-(i)Monte Carlo configuration ${x_{i}}$ for $beta=20,40,80$,respectively
@@ V.2 Action Prediction
With the trained networks it is ready for estimation actions from those at two other temperatures. By pretending having neither the network nor the analytical Lagrangian density at $\beta_{3}$, we start to find the action of an arbitrary sample at at $\beta_{3}$ by making use of two trained neural networks at $\beta_{1}$ and $\beta_{2}$ according to the procedure presented in Section III using Eq.6. As we have already obtained three networks at three temperatures, we will choose any two of them as $\beta_{1}$ and $\beta_{2}$ to predict the remaining one $\beta_{3}$ whose network becomes idle. The predicted total action, kinetic part (proportional to $\beta^{-1}$) and potential part (proportional to $\beta$) of the 2-to-1 prediction and analytical actions are show in the first, second and third column respectively in Figure1. Clearly the first ($\beta_{40,80}$ to $\beta_{20}$) and last ($\beta_{20,40}$ to $\beta_{80}$) rows are extrapolation cases and the second ($\beta_{20,80}$ to $\beta_{40}$) is the interpolation. We can find that the interpolation case, i.e. the $\beta_{20,80}$ to $\beta_{40}$, works best.

### v2

@@ III Neural Networks for Quantum Statistics
where $\Delta\tau=\beta/N_{\tau}$, the first term is part of kinetic term denoted as $K\equiv N_{\tau}\sum(\Delta x)^{3}{(\Delta\Phi)^{2}}$, and the second term includes all the time-independent terms denoted as $V\equiv N_{\tau}^{-1}\sum(\Delta x)^{3}{[(\nabla\Phi)^{2}+V(\Phi)]}$. The temperature dependence of these two terms are different and is separable with regard to the quantum fields. It is evident that once actions of a sample at two given temperatures can be estimated as
@@
$V_{k}(x)=\frac{\lambda_{k}}{4}\left(x^{2}-\frac{\mu_{k}^{2}}{2\lambda_{k}}\right)^{2}.$ (9)
Here $\lambda_{k}$ is the coupling constant with mass dimension $[\lambda_{k}]=[m]^{5}$ and a parameter $\mu_{k}$ with mass dimension $[\mu_{k}]=[m]^{3/2}$. Suppose the particle mass parameter (in the kinetic energy term) is m, we will then work with dimensionless quantities rescaled by proper powers of m (e.g., $\mu^{2}_{k}/m^{3}\rightarrow\mu^{2}_{k}$ and $\lambda_{k}/m^{5}\rightarrow\lambda_{k}$) throughout this paper. The properties of the system at finite temperature $T$ can be described through the partition function
@@ V.1 Action Estimation
In this section we show the training and validation stages of the CANs. It will be seen that the CANs can succeed to reproduce the correct action of most configurations at different temperature without knowing the analytical form of the Lagrangian density (unsupervisely). To train the network we have prepared three ensembles at $\beta=T^{-1}=$20, 40 and 80 with the MCMC simulation. We train the CANs to estimate the probability of each sample in these three ensembles for 10,000 epochs. From Figure 2 we can find the $\beta=40$ case converges fastest, while the $\beta=80$ case is slower and $\beta=80$ slowest. This means the network is good at learning the ensemble with different kinds of configurations, because both the kinetic and potential terms will contribute to the averaged loss equally. Two limit cases is harder to learn because the network will focus on the kinetic/potential part. As a result the loss from the other part will be more difficult to reduce. Nevertheless with a large enough number of epoch(larger than 2000) all the networks has been convergent.
Figure 2: Model loss for $\beta=$20, 40, 80, respectively
@@
Figure 3: (a)-(c) Comparisons of action distribution by MCMC(analytical) and CANs; (d)-(f) Comparisons of action on testing data by MC and CANs; (g)-(i)Monte Carlo configuration ${x_{i}}$ for $\beta=$20, 40, 80, respectively
@@ V.2 Action Prediction
With the trained networks it is ready for estimation actions from those at two other temperatures. By pretending having neither the network nor the analytical Lagrangian density at $\beta_{3}$, we start to find the action of an arbitrary sample at $\beta_{3}$ by making use of two trained neural networks at $\beta_{1}$ and $\beta_{2}$ according to the procedure presented in Section III using Eq.6. As we have already obtained three networks at three temperatures, we will choose any two of them as $\beta_{1}$ and $\beta_{2}$ to predict the remaining one $\beta_{3}$ whose network becomes idle. The predicted total action, kinetic part (proportional to $\beta^{-1}$) and potential part (proportional to $\beta$) of the 2-to-1 prediction and analytical actions are show in the first, second and third column respectively in Figure 4. Clearly the first ($\beta_{40,80}$ to $\beta_{20}$) and last ($\beta_{20,40}$ to $\beta_{80}$) rows are extrapolation cases and the second ($\beta_{20,80}$ to $\beta_{40}$) is the interpolation. In these figures we list two quantities, i.e. $R^{2}$ and $\bar{D}$ to show the predicting ability of the CANS. $R^{2}$ is the mean value of the square of the predicting error $R^{2}=\langle{(S_{True}-S_{CANs})^{2}}\rangle$ and the $\bar{D}$ is the mean of the distance to red line with the slope 1. We can find that the interpolation case, i.e. the $\beta_{20,80}$ to $\beta_{40}$, works best.